OpenStudy (anonymous):

The length of a rectangular garden is 20 feet longer than the width, w. which equation best describes the gardens perimeter , p

OpenStudy (anonymous):

w = w l = 20 + w perimeter = 2l + 2w perimeter = 2w + 2(20+w) perimeter = 2w + 40+ 2w = 4w + 20

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

np glad to help

OpenStudy (anonymous):

can you help with a little more?